# Correlation between the first-reaction time and the acquired boundary local time

### v1

@@ III.3 Regularly packed obstacles in 2d
Figure 4: Mean FRT and correlation coefficient for the domain shown in Fig. 1(a) with $m=5$. Numerical results are obtained by Monte Carlo simulations with $N=10^{6}$ particles that are initially uniformly distributed inside the domain. Symbols correspond to different reactivities: $q=10^{3}$ (squares), $q=10^{2}$ (downward triangles), $q=10^{1}$ (upward triangles), and $q=10^{0}$ (diamonds). (a) Mean FRT as a function of the scaled disk radius $R/R_{\mathrm{max}}$. The solid black curve represents the mean first-passage time (i.e., $q=\infty$), and colored dashed curves show the annulus approximation (44). (b) Correlation coefficient $C_{q}(\circ)$ as a function of scaled disk radius. Colored dashed curves correspond to the annulus approximation (66), in which $L$ is replaced by $L_{\mathrm{eff}}$ from Eq. (43).
@@
The correlation coefficient exhibits a similar trend. The equivalent annulus approximation from Eq. (66) remains accurate for $R/R_{\mathrm{max}}<0.3$ across all selected reactivities. In the high-reactivity limit $q\to\infty$, the power law $C_{q}(\circ)\propto 1/q$ is recovered, as seen from the separation between the blue ($q=10^{3}$) and orange ($q=10^{2}$) dashed curves shown in Fig. 4(b). As expected, when $q\to 0$, one has $C_{q}(\circ)\to 1$. As $R\to R_{\mathrm{max}}$, $C_{q}(\circ)$ shows a sharp decrease.
Figure 5: Role of the obstacle-lattice density $m$ at fixed reactivity $q=10^{2}$. Numerical results obtained by Monte Carlo simulations with $N=10^{6}$ particles that are initially uniformly distributed inside the domain shown in Fig. 1(a). Numerical results are shown for: $m=3$ (squares), $m=5$ (downward triangles), $m=7$ (upward triangles), and $m=9$ (diamonds). (a) Mean FRT as a function of the scaled disk radius $R/R_{\mathrm{max}}$. Colored dashed curves exhibit the annulus approximation (44) for each $m$. (b) Correlation coefficient $C_{q}(\circ)$ as a function of the scaled disk radius. Colored dashed curves exhibit the annulus approximation (66), in which $L$ is replaced by $L_{\mathrm{eff}}$ from Eq. (43).
@@ III.4 Randomly packed obstacles in two dimensions
Figure 6: Numerical results obtained by Monte Carlo simulations with $N=10^{5}$ particles that are initially uniformly distributed inside the domain shown in Fig. 1(b), with different reactivities: $q=10^{2}$ (squares), $q=10^{1}$ (downward triangles), $q=10^{0}$ (upward triangles), and $q=10^{-1}$ (diamonds). (a) Mean FPT as a function of the scaled disk radius $R/R_{\mathrm{max}}$. The black solid line represents the mean first-passage time (i.e., $q\to\infty$), and colored dashed lines exhibit the annulus approximation (44). (b) Correlation coefficient $C_{q}(\circ)$ as a function of the scaled disk radius. Colored dashed lines exhibit the annulus approximation (66), in which $L$ is replaced by $L_{\mathrm{eff}}$ from Eq. (43).
In addition to the regular square-lattice packing, we examine another configuration in which the obstacles are packed randomly. To compare with previous results with $m=5$, we consider the configuration shown in Fig. 1(b) (see Appendix B for coordinates), with one reactive target at the center and 24 obstacles packed randomly. The maximal radius is determined by the 50% area occupation fraction. We consider the mean FRT (Fig. 6(a)) and the correlation coefficient (Fig. 6(b)) as before, where the annulus approximations (44) and (66) are also utilized with the effective radius given by Eq. (43). Compared with the regular packing, the annulus approximation also works when $R/R_{\mathrm{max}}<0.3$ for both mean FRT and correlation coefficient. The main difference arises in the limit $R\to R_{\mathrm{max}}$, in which the mean FRT is finite, and the correlation coefficient remains positive because the target remains accessible from any point in the domain (i.e., there is no geometric isolation of the target).
@@ III.5 Regularly packed obstacles in three dimensions
Figure 7(a) compares the Monte Carlo results at $m=3$ for the mean FRT with this shell approximation (67b) (colored dashed lines). The black solid curve corresponds to the perfect absorption limit $q\to\infty$. Unlike the two-dimensional case, the approximation agrees well with simulations until a larger value of $R/R_{\mathrm{max}}$; especially for $q=10^{-1}$, there is still good agreement even at $R=R_{\mathrm{max}}$. In this limit, all particles can reach the reactive target within a finite FRT, since the three-dimension constraint does not isolate the target like in the two-dimensional case. The correlation coefficient (Fig. 7(b)) exhibits a similar trend. The equivalent shell approximation (72) remains accurate for $R/R_{\mathrm{max}}<0.3$. When the obstacle radius is close to the maximum, $C_{q}$ exhibits a decrease, but remains strictly positive.
Figure 7: Numerical results obtained by Monte Carlo simulations with $N=10^{4}$ particles that are initially uniformly distributed inside the domain shown in Fig. 1(c), with different reactivities: $q=10^{3}$ (squares), $q=10^{2}$ (downward triangles), $q=10^{1}$ (upward triangles), $10^{0}$ (diamonds), and $q=10^{-1}$ (dots). (a) Mean FRT as a function of the scaled ball radius $R/R_{\mathrm{max}}$. The black solid line represents the mean first-passage time (i.e., $q\to\infty$), and colored dashed lines exhibit the shell approximation (67b). (b) Correlation coefficient $C_{q}(\circ)$ as a function of the scaled ball radius. Colored dashed lines exhibit the shell approximation (72), in which $L$ is replaced by $L_{\mathrm{eff}}$ from Eq. (45).
@@ IV Discussion and conclusion
In this study, we have systematically investigated the statistical correlation between two fundamental quantities of diffusion-mediated surface reactions: the first-reaction time $\tau$ and the acquired boundary local time $\ell_{\tau}$. We developed a general theoretical framework for accessing their joint probability distribution and for calculating the correlation coefficient $C_{q}$. We derived explicit analytical expressions for $C_{q}$ in several basic domains, including the ball, the disk, the concentric spherical shell, and the concentric circular annulus. These exact results served as a benchmark for validating our Monte Carlo simulation technique based on the walk-on-spheres and the escape-from-a-layer algorithms. As these correlations are affected by the domain geometry, the analysis was then extended to disordered environments with interior obstacles (2D regular and random packings, as well as 3D regular packing). By employing concentric shell and annulus approximations, we examined the influence of geometric disorder by varying the number of obstacles and their radius. An interesting finding in two-dimensional regular configurations of obstacles is the emergence of a non-monotonic behavior in the mean FRT, a direct consequence of geometrical constraints. In the limiting geometric configurations that completely isolate a particle from the target, mean FRT diverges while the cross-moment vanishes. Across all other geometries, we confirmed the universal limiting behaviors of the correlation: $C_{q}\to 1$ as $q\to 0$, indicating perfect correlation when reactions are rare and boundary exploration is extensive; and $C_{q}\to 0$ as $q\to\infty$, reflecting the decoupling due to instantaneous absorption upon the first arrival. An important theoretical perspective is the analysis of the FRT, the BLT and their correlations for particles diffusing in porous media, in which pores are separated by narrow passages (as in the case of regularly arranged obstacles in two dimensions). Developing refined analytical tools for such poorly connected structures would be valuable.
@@ A.1 Ball
where $\Gamma={\partial\Omega}$, $\alpha=\sqrt{p/D}$, $\mu_{0}^{(p)}=\alpha i_{0}^{\prime}(\alpha R)/i_{0}(\alpha R)$, $V_{0}^{(p)}(\bm{x})=(1/\sqrt{|\Gamma|})i_{0}(\alpha|\bm{x}|)/i_{0}(\alpha R)$, and we used the identity $\int_{\Omega}\mathrm{d}\bm{x}_{0}\ V_{0}^{(p)}(\bm{x}_{0})=D\mu_{0}^{(p)}\int_{\partial\Omega}\mathrm{d}\bm{x}\ V_{0}^{(p)}(\bm{x})/p$. Also, one has

### v2

@@ III.3 Regularly packed obstacles in 2d
(a) Mean FRT
(b) Correlation coefficient
Figure 4: Mean FRT and correlation coefficient for a regular arrangment of obstacles (the domain shown in Fig. 1(a) with $m=5$). Symbols correspond to different reactivities: $q=10^{3}$ (squares), $q=10^{2}$ (downward triangles), $q=10^{1}$ (upward triangles), and $q=10^{0}$ (diamonds). Numerical results are obtained by Monte Carlo simulations with $N=10^{6}$ particles that are initially uniformly distributed inside the domain(for $q=10^{3}$, we used $N=10^{7}$ and $\varepsilon=10^{-4}$). (a) Mean FRT as a function of the scaled disk radius $R/R_{\mathrm{max}}$. The solid black curve represents the mean first-passage time (i.e., $q=\infty$), and colored dashed curves show the annulus approximation (46). (b) Correlation coefficient $C_{q}(\circ)$ as a function of scaled disk radius. Colored dashed curves correspond to the annulus approximation (68), in which $L$ is replaced by $L_{\mathrm{eff}}$ from Eq. (45). Occupation fraction of obstacles is shown on the horizontal top axis.
@@
It is worth noting that the annulus approximation was introduced solely to illustrate the effect of the excluded volume in diluted configurations of obstacles. This basic approximation is not expected to control connectivity and narrow-passage effects that dominate first-passage and reaction statistics, as witnessed by Fig. 4 (see deviations at large $R$). In this light, it was unexpected that such a crude approximation was good enough to accurately reproduce a significant decrease of the mean FRT time as $R/R_{\rm max}$ increases up to $0.4$ (that corresponds to the occupation fraction of around $10\%$).
The correlation coefficient exhibits a similar trend. The equivalent annulus approximation from Eq. (68) remains accurate for $R/R_{\mathrm{max}}<0.3$ across all selected reactivities. In the high-reactivity limit $q\to\infty$, the power law $C_{q}(\circ)\propto 1/q$ is recovered, as seen from the separation between the blue ($q=10^{3}$) and orange ($q=10^{2}$) dashed curves shown in Fig. 4(b). As expected, when $q\to 0$, one has $C_{q}(\circ)\to 1$. In the limit $R\to R_{\mathrm{max}}$, $C_{q}(\circ)$ shows a sharp decrease.
(a) Mean FRT
(b) Correlation coefficient
Figure 5: Role of the obstacle-lattice density $m$ at fixed reactivity $q=10^{2}$. Numerical results are shown for: $m=3$ (squares), $m=5$ (downward triangles), $m=7$ (upward triangles), and $m=9$ (diamonds). We used Monte Carlo simulations with $N=10^{6}$ particles that are initially uniformly distributed inside the domain shown in Fig. 1(a). (a) Mean FRT as a function of the scaled disk radius $R/R_{\mathrm{max}}$. Colored dashed curves show the annulus approximation (46) for each $m$. (b) Correlation coefficient $C_{q}(\circ)$ as a function of the scaled disk radius. Colored dashed curves show the annulus approximation (68), in which $L$ is replaced by $L_{\mathrm{eff}}$ from Eq. (45).
@@ III.4 Randomly packed obstacles in two dimensions
(a) Mean FRT
(b) Correlation coefficient
Figure 6: Mean FRT and correlation coefficient for a random arrangment of obstacles (the domain shown in Fig. 1(b)). Symbols correspond to different reactivities: $q=10^{2}$ (squares), $q=10^{1}$ (downward triangles), $q=10^{0}$ (upward triangles), and $q=10^{-1}$ (diamonds). Numerical results obtained by Monte Carlo simulations with $N=10^{6}$ particles that are initially uniformly distributed inside the domain. (a) Mean FRT as a function of the scaled disk radius $R/R_{\mathrm{max}}$. The black solid line represents the mean first-passage time (i.e., $q\to\infty$), and colored dashed lines show the annulus approximation (46). (b) Correlation coefficient $C_{q}(\circ)$ as a function of the scaled disk radius. Colored dashed lines show the annulus approximation (68), in which $L$ is replaced by $L_{\mathrm{eff}}$ from Eq. (45).
In addition to the regular square-lattice packing, we examine another configuration in which the obstacles are packed randomly. To compare with previous results for $m=5$, we consider the configuration shown in Fig. 1(b) (see Appendix B for coordinates), with one reactive target at the center and 24 obstacles packed randomly. The maximal radius is determined by the 50% area occupation fraction. We consider the mean FRT (Fig. 6(a)) and the correlation coefficient (Fig. 6(b)) as before, where the annulus approximations (46) and (68) are also utilized with the effective radius given by Eq. (45). Compared with the regular packing, the annulus approximation also works when $R/R_{\mathrm{max}}<0.3$ for both mean FRT and correlation coefficient. The main difference arises in the limit $R\to R_{\mathrm{max}}$, in which the mean FRT is finite, and the correlation coefficient remains positive because the target remains accessible from any point in the domain (i.e., there is no geometric isolation of the target).
@@ III.5 Regularly packed obstacles in three dimensions
Figure 7(a) compares the Monte Carlo results at $m=3$ for the mean FRT with this shell approximation (69b) (colored dashed lines). The black solid curve corresponds to the perfect absorption limit $q\to\infty$. Unlike the two-dimensional case, the approximation agrees well with simulations until a larger value of $R/R_{\mathrm{max}}$; for instance, for $q=10^{-1}$, there is still good agreement even at $R=R_{\mathrm{max}}$. In this setting, all particles can reach the reactive target within a finite FRT, since the three-dimension constraint does not isolate the target like in the two-dimensional case. The correlation coefficient (Fig. 7(b)) exhibits a similar trend. The equivalent shell approximation (74) remains accurate for $R/R_{\mathrm{max}}<0.3$. When the obstacle radius is close to the maximum, $C_{q}$ exhibits a decrease, but remains strictly positive.
(a) Mean FRT
(b) Correlation coefficient
Figure 7: Mean FRT and correlation coefficient for a regular arrangment of obstacles inside the ball (the domain shown in Fig. 1(c)). Symbols correspond to different reactivities: $q=10^{3}$ (squares), $q=10^{2}$ (downward triangles), $q=10^{1}$ (upward triangles), $10^{0}$ (diamonds), and $q=10^{-1}$ (dots). Numerical results obtained by Monte Carlo simulations with $N=10^{4}$ particles that are initially uniformly distributed inside the domain. (a) Mean FRT as a function of the scaled ball radius $R/R_{\mathrm{max}}$. The black solid line represents the mean first-passage time (i.e., $q\to\infty$), and colored dashed lines show the shell approximation (69b). (b) Correlation coefficient $C_{q}(\circ)$ as a function of the scaled ball radius. Colored dashed lines show the shell approximation (74), in which $L$ is replaced by $L_{\mathrm{eff}}$ from Eq. (47).
@@ IV Discussion and conclusion
In this study, we have systematically investigated the statistical correlation between two fundamental quantities of diffusion-mediated surface reactions: the first-reaction time $\tau$ and the acquired boundary local time $\ell_{\tau}$. We developed a general theoretical framework for accessing their joint probability density ${\mathcal{P}}(\ell,t|\bm{x}_{0})$. For a broad class of diffusion-controlled reactions, Eq. (8) decouples the effects of chemical kinetics (incorporated through the probability density $\psi(\ell)$ of the imposed threshold $\hat{\ell}$) and diffusive dynamics (represented via the probability density $U(\ell,t|\bm{x}_{0})$ of the first-crossing time). This multiplicative form is a consequence of the assumed independence of these two “ingredients”. Despite an apparent simplicity of this representation, the considered model is fairly general; for instance, it includes the most common scenario of a partially reactive region with a constant reactivity as a special case. The derived representation of the joint probability density sheds light on correlations between $\tau$ and $\ell_{\tau}$ that are characterized through the correlation coefficient $C_{q}$. We derived explicit analytical expressions for $C_{q}$ in several basic domains, including the ball, the disk, the concentric spherical shell, and the concentric circular annulus. These exact results served as a benchmark for validating our Monte Carlo simulation technique based on the walk-on-spheres and the escape-from-a-layer algorithms. As these correlations are affected by the domain geometry, the analysis was then extended to disordered environments with interior obstacles (2D regular and random packings, as well as 3D regular packing). By employing concentric shell and annulus approximations, we examined the influence of geometric disorder by varying the number of obstacles and their radius. An interesting finding in two-dimensional regular configurations of obstacles is the emergence of a non-monotonic behavior in the mean FRT, a direct consequence of geometrical constraints. In the limiting geometric configurations that completely isolate a particle from the target, mean FRT diverges while the cross-moment vanishes. Across all other geometries, we confirmed the universal limiting behaviors of the correlation: $C_{q}\to 1$ as $q\to 0$, indicating perfect correlation when reactions are rare and boundary exploration is extensive; and $C_{q}\to 0$ as $q\to\infty$, reflecting the decoupling due to instantaneous absorption upon the first arrival. An important theoretical perspective is the analysis of the FRT, the BLT and their correlations for particles diffusing in porous media, in which pores are separated by narrow passages (as in the case of regularly arranged obstacles in two dimensions). Developing refined analytical tools for such poorly connected structures would be valuable.
@@ A.1 Ball
where $\Gamma={\partial\Omega}$, $\alpha=\sqrt{p/D}$, $\mu_{0}^{(p)}=\alpha i_{0}^{\prime}(\alpha R)/i_{0}(\alpha R)$, $V_{0}^{(p)}(\bm{x})=(1/\sqrt{|\Gamma|})i_{0}(\alpha|\bm{x}|)/i_{0}(\alpha R)$, and we used the identity
$\int_{\Omega}\mathrm{d}\bm{x}_{0}\ V_{0}^{(p)}(\bm{x}_{0})=\frac{D\mu_{0}^{(p)}}{p}\int_{\partial\Omega}\mathrm{d}\bm{x}\ V_{0}^{(p)}(\bm{x}).$
Also, one has